If a string of length l is initially at rest in equilibrium position and each of its points is given the velocity dy/dt= b sin^3πx/l find the displacement





1
Expert's answer
2022-01-04T13:42:28-0500

The displacement y(x,t) is given by the equation:

"y_{tt}=a^2y_{xx}"

boundary conditions:

"y(0,t)=y(l,t)=0"

"y_t(0)=b sin(3\u03c0x\/l)"

"y(x,0)=0"


solution:

"y(x,t)=Bsin(n\pi x/l)(Ccos(n\pi at/l)+Dsin(n\pi at/l))"

then:

"y(x,0)=Bsin(n\pi x/l)C=0 \implies C=0"

"y(x,t)=Bsin(n\pi x/l)Dsin(n\pi at/l)"

the most general solution:

"y(x,t)=\sum B_nsin(n\pi x/l)sin(n\pi at/l)"


then:

"y_t=\sum B_n(n\pi a/l)sin(n\pi x/l)sin(n\pi at/l)"


"y_t(0)=b sin(3πx/l)=\sum B_n(n\pi a/l)sin(n\pi x/l)sin(n\pi at/l)"


"b=B_3(3\pi a/l),B_1=B_2=B_4=...=0"


"B_3=\frac{bl}{3\pi a}"


"y(x,t)=B_3sin(3\pi x/l)sin(3\pi at/l)=\frac{bl}{3\pi a}sin(3\pi x/l)sin(3\pi at/l)"